The comparative financial statements prepared at December 31 for Pinnacle Plus showed the following summarized data: Current1-b. Are the current year results better, or worse, than those for the previous year? Better Worse 2-a. Compute the net profi4-a. Stockholders equity totaled $103,000 at the beginning of the previous year. Compute the return on equity (ROE) ratios f6-b. Is debt providing financing for a larger or smaller proportion of the companys asset growth? Larger Proportion Smaller
Answer #1

1a. Gross profit ratio = Gross profit / Sales * 100

Current year = $67,000 / 125,000 * 100 = 53.6%

Previous year = $57,000 / 111,000 * 100 = 51.4%

1b. Better

2a. Net profit margin = Net profit / Sales * 100

Current year = $16,590 / 125,000 *100 = 13.3%

Previous year = $12,110 / 111,000 *100 = 10.9%

2b. Better

3a. Earning per share = Net income / Weighted average outstanding common stock

Current year = $16,590 / 9,180 = $1.81

Previous year = $12,110 / 9,180 = $1.32

3b. Better

4a. Return on equity ratio = Net income / Average stockholders equity

Current year = $16,590 / (130,340+120,650/2)

Current year = $16,590 / $125,495 * 100 = 13.2%

Previous year = $12,110 / (120,650+103,000/2)

Previous year = $12,110 / $111,825 * 100 = 10.8%

4b. Better

5a. Fixed asset turnover ratio = Sales revenue / Average fixed assets

Current year = $125,000 / (98,000+108,000/2)

Current year = $125,000 / $103,000 = 1.21

Previous year = $111,000 / (108,000+113,000/2)

Previous year = $111,000 / $110,500 = 1.00

5b. Better

6a. Debt to asset ratio = Total debt / Total assets

Current year = $89,075 / 219,415 = 0.41

Previous year = $78,350 / 199,000 = 0.39

6b. Larger proportion

7a. Times interest earned ratio = Income before interest and tax / interest expense

Current year = $67,000 - 39,000 / 4,300 = 6.5

Previous year = $57,000 - 35,400 / 4,300 = 5

7b. Better

8a. Price earning ratio = Market price per share / earning per share

Current year = $21 / 1.81 = 11.6

Previous year = $18 / 1.32 = 13.6

8b. Less optimistic